OpenStudy (anonymous):

How many solutions does the equation have? 4(2x=3) = 2(3x-4)

Directrix (directrix):

4(2x=3) = 2(3x-4) -------------- This has two equality symbols. Should it be: 4(2x-3) = 2(3x-4) or something else?

OpenStudy (anonymous):

A. one solution B. infinite solutions C. no solution these are the possible answers

OpenStudy (anonymous):

no it has 1 hahaaha

OpenStudy (anonymous):

solve the equation in the end you end with a single x

OpenStudy (anonymous):

thanks

Directrix (directrix):

@Bryan11 How did you get one solution from: 4(2x=3) = 2(3x-4) ?

OpenStudy (anonymous):

hey that's clearly not an equal sign @Directrix

OpenStudy (anonymous):

hahaha @Directrix